OpenStudy (anonymous):

Write the slope-intercept form of an equation of a line parallel to y = -2x - 9 and passing through the point (-2,3).

OpenStudy (anonymous):

answer please??

OpenStudy (anonymous):

y=mx+b m=slope b= y-intercept y = -2x - 9 What is your slope in this case?

OpenStudy (anonymous):

To find an eqn of a line: all you need is a point (-2,3) and a slope. Plug them in to y=mx+b to find b, the y-intercept. Now you'll have your slope and y-intercept. Plug only those into y=mx+b, and there you have it. Your parallel line.